# The Fundamental Limit of Jet Tagging

### v1

@@ Abstract
We fix the size of the classifier training dataset to 1M boosted top quark jets and 1M QCD jets and compare different neural network architectures, as was done in Ref. [3], only without knowing the optimal solution. Since that original work, there have been many innovations in jet tagging and we consider a selection of state-of-the-art and/or widely used models including Deep Sets [28] without IRC safety (Particle Flow Networks [29]) and with attention (Deep Sets with Attention [27]); Lorentz-equivariant models such as LorentzNet [30] and Pelican [31]; transformers such as ParT [6] and OmniLearn (trained from scratch) [32] and a classifier based on the transformer architecture of Ref. [17], where we replace the output layer with a binary cross entropy layer (Baseline Transformer); and the fine-tuned OmniLearn foundation model [32]. In all cases, we have used the settings described in the corresponding papers without additional hyperparameter optimization. Our results are displayed in Fig. 1. All taggers show comparable performance, which is also similar to the performance achieved on the datasets of Refs. [3, 6]. However, even the most powerful current taggers fall well short of optimal performance; for example, with a top-tagging efficiency of 0.5, the taggers exhibit a background rejection of at most about 600, whereas optimal performance corresponds to a background rejection of about 3000.
Figure 1: ROC curves for top quark versus QCD jet classification. The performance of the likelihood ratio classifier is labeled ’Optimal’ and the other curves correspond to different classifiers trained with 1M jets of each class. The parentheses in the legend indicate the rejection (inverse false positive rate) at a top quark efficiency of 50% and the corresponding error from training 5 classifiers independently.
To analyze our results in more depth, we first investigate whether the performance of the taggers can be further improved by training on an even larger dataset. To do this, we compare the optimal performance with the ROC curves for a selected tagger trained on training datasets of different sizes, from 100 to 10M events. The results are shown in Fig. 2 for the classifier based on the transformer architecture of Ref. [17] (Baseline Transformer in Fig. 1). Of course, the performance of the tagger initially improves with the size of the training dataset, but from about 1M training events a saturation becomes apparent. This means that even with significantly larger training data sets, optimal performance cannot be expected.
@@
Figure 3: ROC curves for top quark versus QCD jet classification for decreasing number of jet constituents $N_{c}$. The performance of the likelihood ratio classifier is shown as solid lines, while the dashed lines correspond to the baseline transformer classifier. The parentheses in the legend indicate the rejection (inverse false positive rate) at a top quark efficiency of 50%.

### v2

@@ Abstract
We fix the size of the classifier training dataset to 10M boosted top quark jets and 10M QCD jets and compare different neural network architectures, similar to the widely-used benchmark comparison of Ref. [3], only now we know the optimal solution. Since that original work, there have been many innovations in jet tagging and we consider a selection of state-of-the-art and/or widely used models including ParticleNet [5], Deep Sets [28] without IRC safety (Particle Flow Networks [29]) and with attention (Deep Sets with Attention [27]); Lorentz-equivariant models such as LorentzNet [30] and Pelican [31]; transformers such as ParT [6] and OmniLearn (trained from scratch) [32] and a classifier based on the transformer architecture of Ref. [17], where we replace the output layer with a binary cross entropy layer (Baseline Transformer); and the fine-tuned OmniLearn foundation model [32]. In all cases, we have used the settings described in the corresponding papers without additional hyperparameter optimization. Our results are displayed in Fig. 1. All taggers show comparable performance, which is also similar to the performance achieved on the datasets of Refs. [3, 6]. However, even the most powerful current taggers fall well short of optimal performance; for example, with a top-tagging efficiency of 0.5, the taggers exhibit a background rejection of at most about 550, whereas optimal performance corresponds to a background rejection of about 3300.
Figure 1: ROC curves for top quark versus QCD jet classification. The performance of the likelihood ratio classifier is labeled ’Optimal’ and the other curves correspond to different classifiers trained with 10M jets of each class. The parentheses in the legend indicate the rejection (inverse false positive rate) at a top quark efficiency of 50% and the corresponding error from training 5 classifiers independently.
To analyze our results in more depth, we first investigate how the performance of the taggers varies with the size of the training dataset. To do this, we compare the optimal performance with the ROC curves for a selected tagger trained on training datasets of different sizes, from 100 to 10M events. The results are shown in Fig. 2 for the classifier based on the transformer architecture of Ref. [17] (Baseline Transformer in Fig. 1). Of course, the performance of the tagger initially improves with the size of the training dataset, but the improvement from 1M to 10M training events is rather limited. This means that even with significantly larger training data sets, optimal performance cannot be expected (for fixed architecture).
@@
Figure 3: ROC curves for top quark versus QCD jet classification for decreasing number of jet constituents $N_{c}$. The performance of the likelihood ratio classifier is shown as solid lines, while the dashed lines correspond to the baseline transformer classifier. The parentheses in the legend indicate the rejection (inverse false positive rate) at a top quark efficiency of 50%. Note that the dashed lines in this figure correspond to a baseline transformer trained on 1M events.